OpenStudy (anonymous):

18 is 16% of what number I got 30, correct?

OpenStudy (anonymous):

\[\frac{ 18 }{ x}=\frac{ 60 }{ 100 } \]

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

Is this ewuation right? Means-extremes product? (cross cancel)

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok, how so?

OpenStudy (kittiwitti1):

16%? You put 60%, lol

OpenStudy (anonymous):

oops 16

OpenStudy (kittiwitti1):

No no, the proportion says 60/100. That's where you messed up

OpenStudy (anonymous):

wait nvm it's 60% lmao

OpenStudy (kittiwitti1):

:|

OpenStudy (anonymous):

anyways, I'm correct?

OpenStudy (kittiwitti1):

┬─┬ ノ( ゜-゜ノ) (╯°□°)╯︵ ┻━┻

OpenStudy (kittiwitti1):

Yes you got it right (-‸ლ)

OpenStudy (anonymous):

oh, ok thanks.

OpenStudy (kittiwitti1):

no problem ○ x ○

OpenStudy (anonymous):

18 is 16% of what number translate directly in to math 18 means 18 is means equals 16% means \(0.16\) of means \(\times \) what number is a variable, i pick \(n\) equation is \[18=0.16\times n\]

OpenStudy (anonymous):

solve for \(n\) via \[16\div 0.16=n\]

OpenStudy (anonymous):

it is definitely not 30!

OpenStudy (kittiwitti1):

@satellite73 he said he made a mistake, the problem is asking for 60% lol

OpenStudy (anonymous):

she*

OpenStudy (kittiwitti1):

o_e oops sorry